OpenStudy (anonymous):

steps to solve √8a^3?

OpenStudy (anonymous):

2asqrt2a

OpenStudy (anonymous):

\[\sqrt{2^3a^3}\]

OpenStudy (anonymous):

\[\sqrt{2^2a^2 \times 2a}\] \[2a \sqrt{2a}\]